OpenStudy (anonymous):

Let f(x)=8/x. Then the expression f(x+h)−f(x)/h can be written in the form A/(x+h), where A is a constant and A= Using your answer from above we have: limh→0f(x+h)−f(x)/h=

OpenStudy (anonymous):

A=????

hero (hero):

\[f(x) = \frac{8}{x}\]\[f(x+h) = \frac{8}{x + h}\] \(f(x+h) - f(x) = \dfrac{8}{x + h} - \dfrac{8}{x}\) \( = \dfrac{8x}{x(x + h)} - \dfrac{8(x+h)}{x(x+h)}\) \(=\dfrac{8x - 8(x+h)}{x(x+h)}\) \(=\dfrac{8x - 8x -8h}{x(x + h)}\) \(=\dfrac{-8h}{x(x+h)}\) \(\dfrac{f(x+h) - f(x)}{h} = \dfrac{\frac{-8h}{x(x+h)}}{h}\) \(=\dfrac{-8h}{x(x+h)} \div h\) \(=\dfrac{-8h}{x(x+h)} \times \dfrac{1}{h}\) \(=\dfrac{-8}{x+h}\) Thus \(A = -8\) And \(\lim_{h \to 0} \dfrac{f(x+h) - f(x)}{h} = \dfrac{-8}{x + 0} = -\dfrac{8}{x}\)